# High-Dimensional Sparse Clustering via Iterative Semidefinite Programming Relaxed K-Means

### v1

@@ 3.2.1 Innovated Scalable Efficient Estimation (ISEE) subroutine
The ISEE procedure is motivated by two key observations. First, the transformed vector $\tilde{\mathbf{X}}_{i}=\boldsymbol{\Omega}^{\ast}\mathbf{X}_{i}\in\mathbb{R}^{p}$ remains Gaussian, with $\operatorname{Cov}(\tilde{\mathbf{X}}_{i})=\boldsymbol{\Omega}^{\ast}\boldsymbol{\Sigma}^{\ast}\boldsymbol{\Omega}^{\ast}=\boldsymbol{\Omega}^{\ast}.$ Without loss of generality, assume $p$ is an even integer. Consider a subset $A\subset\{1,\ldots,p\}$ of size 2. Assume $i\in G_{1}^{\ast}$, and consider the following decomposition of $\tilde{\mathbf{X}}_{A,i}$ as the sum of deterministic mean part and mean-zero random noise vector:
@@
[…]
where $\mathbf{E}_{A,i}\sim\mathcal{N}(0,(\boldsymbol{\Omega}^{\ast}_{A,A})^{-1})$. Leveraging this relationship, the quantities in (9) can be further expressed as follows (Fan and Lv,, 2016):
@@
Appendix F.2 provides detailed derivations for (10) and (11). Since $p-|A|$ can be comparable or larger than $n$, one may want to leverage the existing regularization method for fitting the regression model (10) to obtain sample estimates of $\boldsymbol{\Omega}^{\ast}_{A,A}$, $\mathbf{E}_{A,i}$ and $(\boldsymbol{\alpha}_{1})_{A}$, which further yields plug-in estimate of $\tilde{\mathbf{X}}_{A,i}$. This process can be repeated for all $A\in\mathcal{A}$ with $\mathcal{A}$ a partition set of $\{1,\cdots,p\}$ consisting of sets of size 2. By combining estimates for $\tilde{\mathbf{X}}_{A,i}$ for all $A\in\mathcal{A}$, we can form an estimate of the full matrix $\tilde{\mathbf{X}}_{G_{1}^{*}}\in\mathbb{R}^{|G_{1}^{*}|\times p}$.
@@
Two issues remain to be addressed. First, the true cluster labels are unknown. Second, we need an implementable ISEE subroutine. Since our algorithm is iterative by nature, we use the cluster assignments in the current iteration as the proxy for true cluster labels to implement the ISEE subroutine, and then iteratively update the estimates for $\tilde{\mathbf{X}}$ and cluster assignments. We next explain the ISEE subroutine for a given pair of cluster assignment $(G_{1},G_{2})$.
ISEE subroutine.
[…]
3.2.2 Full algorithm with iterative variable selection via ISEE
@@
Given the ISEE subroutine, we extend Algorithm 2 to the unknown covariance scenario. In the variable selection step, we threshold $\hat{\tilde{\boldsymbol{\mu}}}_{1}^{t}-\hat{\tilde{\boldsymbol{\mu}}}_{2}^{t}$, instead of $\bar{\tilde{\mathbf{X}}}_{\hat{G}_{1}^{t-1}}-\bar{\tilde{\mathbf{X}}}_{\hat{G}_{2}^{t-1}}$. In the clustering step, we replace $\tilde{\mathbf{X}}_{\hat{S}^{t},\cdot}$ with the estimated transformed data matrix $\hat{\tilde{\mathbf{X}}}^{t}_{\hat{S}^{t},\cdot}$. We replace $\boldsymbol{\Sigma}_{\hat{S}^{t},\hat{S}^{t}}$ with the pooled sample covariance, combining the sample covariances of $\mathbf{X}_{\hat{S}^{t},\hat{G}^{t-1}_{1}}$ and $\mathbf{X}_{\hat{S}^{t},\hat{G}^{t-1}_{2}}$. We leverage the convergence rate of the ISEE procedure (Fan and Lv,, 2016) to guide the choice of threshold for $\hat{\tilde{\boldsymbol{\mu}}}_{1}^{t}-\hat{\tilde{\boldsymbol{\mu}}}_{2}^{t}$. Specifically, for each $k\in\{1,2\}$ and for each block $A_{\ell}$ with $\ell\in\{1,\ldots,L\}$, under some regularity conditions (Fan and Lv,, 2016), we have:
$\bigl{\|}\bigl{(}(\hat{\tilde{\boldsymbol{\mu}}}_{1})_{A_{\ell}}-$ $(\hat{\tilde{\boldsymbol{\mu}}}_{2})_{A_{\ell}}\bigr{)}-\bigl{(}({\tilde{\boldsymbol{\mu}}}_{1}^{\ast})_{A_{\ell}}-({\tilde{\boldsymbol{\mu}}}_{2}^{\ast})_{A_{\ell}}\bigr{)}\bigr{\|}_{2}\lesssim C_{6}\max\{J\log p/n,\sqrt{\log p/n}\},$ (12)
@@
where $J$ is the maximum number of nonzero off-diagonal entries per row of $\boldsymbol{\Omega}^{\ast}$, and $C_{6}=\|\boldsymbol{\Omega}^{\ast}_{A_{\ell},A_{\ell}}\|_{2}+\|(\hat{\boldsymbol{\alpha}}_{1})_{A_{\ell}}\|_{2}+\|(\hat{\boldsymbol{\alpha}}_{2})_{A_{\ell}}\|_{2}$. This bound comes from the convergence rates of $\|\hat{\boldsymbol{\Omega}}_{A_{\ell},A_{\ell}}-\boldsymbol{\Omega}^{\ast}_{A_{\ell},A_{\ell}}\|_{2}$ and $\|(\hat{\boldsymbol{\alpha}}_{k})_{A_{\ell}}-(\boldsymbol{\alpha}_{k})_{A_{\ell}}\|_{2}$ proved in Fan and Lv, (2016). The detailed derivation is provided in Appendix F.2. In the standard sparse high-dimensional regime of $J\sqrt{\log p/n}=o(1)$, the term $\sqrt{\log p/n}$ dominates the bound. Since ISEE only involves small submatrices of size $2\times 2$ or $3\times 3$, the terms in $C_{6}$ are bounded and contribute only constant factors. To account for these constants and suppress false discoveries, we include a slowly diverging factor $\sqrt{\log n}$ in our threshold. Combining all the pieces, we set the variable selection threshold as $\sqrt{\log n\log p/n}$. The final algorithm incorporating these thresholding principles is presented in Algorithm 3.
@@
2:$\hat{G}_{1}^{0},\hat{G}_{2}^{0}\leftarrow$ cluster estimates obtained by running spectral clustering on $\mathbf{X}$triangleright Initialization
@@
4:     $\hat{\tilde{\boldsymbol{\mu}}}_{1}^{t},\hat{\tilde{\boldsymbol{\mu}}}_{2}^{t},\hat{\tilde{\mathbf{X}}}^{t},\leftarrow\texttt{ISEE}(\mathbf{X},\hat{G}_{1}^{t-1},\hat{G}_{2}^{t-1})$$\triangleright$ Estimated transformation (Algorithm 4)
5:     $\hat{S}^{t}\leftarrow\left\{j\in\{1,\ldots,p\}:|(\widehat{\tilde{\boldsymbol{\mu}}}_{1}^{t}-\widehat{\tilde{\boldsymbol{\mu}}}_{2}^{t})_{j}|>\sqrt{\log n\log p/n}\right\}$triangleright Variable selection
@@
7:     $\hat{G}_{1}^{t},\hat{G}_{2}^{t}\leftarrow\texttt{SDPkmeans}(\hat{\tilde{\mathbf{X}}}_{\hat{S}^{t},\cdot},\hat{\boldsymbol{\Sigma}}_{\hat{S}^{t},\hat{S}^{t}})$triangleright Clustering (Algorithm 1)
@@
4 Numerical results
For all simulations, we consider a symmetric bicluster Gaussian mixture model with equal cluster sizes and a common covariance. For data dimension $p$ and sample size $n$, the data are generated as
@@ 4 Numerical results
$\mathbf{X}_{i}\overset{i.i.d.}{\sim}\mathcal{N}(\boldsymbol{\mu}^{\ast},\boldsymbol{\Sigma}^{\ast}),~{}i=1,\dots,n/2,~{}\text{and}~{}\mathbf{X}_{i}\overset{i.i.d.}{\sim}\mathcal{N}(-\boldsymbol{\mu}^{\ast},\boldsymbol{\Sigma}^{\ast}),i=n/2+1,\dots,n.$
@@
We define $\boldsymbol{\beta}^{\ast}:=\boldsymbol{\Omega}^{\ast}\boldsymbol{\mu}^{\ast}$ here and set its support to $S_{0}=[10]$. All non-zero entries of $\boldsymbol{\beta}^{\ast}$ have the same magnitude, chosen to keep the signal strength, measured by the Mahalanobis distance $\Delta_{\boldsymbol{\Sigma}^{\ast}}=2\sqrt{(\boldsymbol{\mu}^{\ast})^{\top}\boldsymbol{\Omega}^{\ast}\boldsymbol{\mu}^{\ast}}=\sqrt{(\boldsymbol{\beta}^{\ast})^{\top}\boldsymbol{\Sigma}^{\ast}\boldsymbol{\beta}^{\ast}},$ as constant, while increasing the noise level through higher ambient dimension $p$. We evaluate performance using the mis-clustering rate, averaged over 200 independent replications. For initialization, we use spectral clustering (Han et al.,, 2023). We set the number of iteration as $T=100$, but used early stopping criterion. The details on this criterion, code implementation, and details required for replicating our results are provided in Appendix G.1.
Table 1: Summary of Simulation Settings. $i,j=1,\ldots,p$.
Methods Evaluated Covariance structure $p$ $n$ Separation
Algorithm 2 (for known $\boldsymbol{\Sigma}^{\ast}$) known, $\boldsymbol{{\Sigma}}^{\ast}=\mathbf{I}_{p}$ 50-5000 200 $\Delta^{2}=4^{2},5^{2}$
Algorithm 3 (for unknown $\boldsymbol{\Sigma}^{\ast}$) unknown, $\Omega^{*}_{i,i+1}=\Omega^{*}_{i+1,i}=\rho$, $\Omega^{*}_{i,i}=1$, $\Omega^{*}_{i,j}=0$ o.w., $\rho\in\{0.2,0.45\}$ 50-400 500 $\Delta_{\boldsymbol{\Sigma}^{\ast}}^{2}=3^{2},4^{2}$
Scenario 1: known isotropic covariance
@@
We evaluate Algorithm 2 under the setting of line 1 of Table 1. We first compare non-sparse clustering methods, including spectral clustering (Han et al.,, 2023), hierarchical clustering, and SDP-relaxed K-means with all features (Chen and Yang,, 2021), with Algorithm 2 initialized by each of them. Here, initialization refers to a rough cluster assignment $\hat{G}_{1}^{0},\hat{G}_{2}^{0}$ used to guide the first variable selection step. Results in Figures 1-(a) and 1-(b) show that while the non-sparse methods fail to adapt to sparsity, Algorithm 2, when properly initialized, is robust to high-dimensional noise and effectively adapts to sparsity. Even moderately good initializations are sufficient, as evidenced by the improved performance of Algorithm 2 when initialized with spectral or SDP K-means clustering. In contrast, poor initializations can hinder recovery, as reflected in the weaker performance of Algorithm 2 initialized by hierarchical clustering. Next, we benchmark our method against existing sparse clustering approaches, including a two-step method (IF-PCA; Jin and Wang,, 2016) and iterative methods such as SAS (Arias-Castro and Pu,, 2017), Sparse K-means (Witten and Tibshirani,, 2010), and CHIME (Cai et al.,, 2019). The results in Figures 1-(c) and 1-(d) highlight the competitiveness of our approach. Details of the baseline methods are provided in Appendix G.2.
@@
Figure 1: Mis-clustering rates of Algorithm 2 compared to sparse and non-sparse baselines. The dotted lines (non-sparse methods) can also be interpreted as the mis-clustering rates of Algorithm 2 immediately after initialization, before any iteration. The settings are provided in line 1 of Table 1.
Scenario 2: Unknown covariance.
We evaluate the performance of Algorithm 3 in the setting with a chain-graph sparse precision matrix with conditional correlation $\rho\in\{0.2,0.4\}$, following the setup described in line 2 of Table 1. Figures 2-(a) to 2-(d) compare Algorithm 3 with two alternative variants for the variable selection step: Algorithm 5 estimates the full precision matrix using graphical lasso (Friedman et al.,, 2008), while Algorithm 6 applies thresholding to the empirical cluster means of the transformed data, using a cutoff derived from the Gaussian maximal inequality. Details of these algorithms are provided in Appendix G.2. The results show that Algorithm 3 outperforms both alternatives, highlighting the benefits of bypassing full precision matrix estimation and employing convergence-rate-guided thresholding. We also compare our method against other sparsity-aware clustering approaches. The result presented in Figures 2-(e) to 2-(h) show that our method outperforms the others.
@@
Figure 2: Mis-clustering rates of Algorithm 3, compared to its variants (Algorithms 5 and 6; plots a–d) and other sparsity-aware baseline methods (plots e-h), under unknown covariance scenario with varying conditional correlations. The settings are provided in line 2 of Table 1.

### v2

@@ 3.2.1 Innovated Scalable Efficient Estimation (ISEE) Subroutine
The ISEE procedure is motivated by two key observations. First, the transformed vector $\tilde{\mathbf{X}}_{i}=\boldsymbol{\Omega}^{\ast}\mathbf{X}_{i}\in\mathbb{R}^{p}$ remains Gaussian, with $\operatorname{Cov}(\tilde{\mathbf{X}}_{i})=\boldsymbol{\Omega}^{\ast}\boldsymbol{\Sigma}^{\ast}\boldsymbol{\Omega}^{\ast}=\boldsymbol{\Omega}^{\ast}.$ Without loss of generality, assume $p$ is an even integer, consider a subset $A\subset[p]$ of size 2, and assume $i\in G_{1}^{\ast}$. We consider the following decomposition of $\tilde{\mathbf{X}}_{A,i}$ as the sum of deterministic mean part and mean-zero random noise vector:
@@
$\hskip-27.0pt\underbrace{\mathbf{X}_{A,i}}_{\hskip 9.01396pt\text{response}\in\mathbb{R}^{|A|}}\hskip-16.00008pt=\hskip 6.99997pt\underbrace{(\boldsymbol{\mu}_{1}^{\ast})_{A}+(\boldsymbol{\Omega}^{\ast}_{A,A})^{-1}\boldsymbol{\Omega}^{\ast}_{A,A^{c}}(\boldsymbol{\mu}_{1}^{\ast})_{A^{c}}}_{:=(\boldsymbol{\alpha}_{1})_{A}\in\mathbb{R}^{|A|}~\text{(intercept)}}~-~\underbrace{(\boldsymbol{\Omega}^{\ast}_{A,A})^{-1}\boldsymbol{\Omega}^{\ast}_{A,A^{c}}}_{\hskip-21.30568pt\text{slope}\in\mathbb{R}^{|A|\times(p-|A|)}}\hskip-16.00008pt\underbrace{\mathbf{X}_{A^{c},i}}_{\hskip 9.01396pt\text{predictor}\in\mathbb{R}^{p-|A|}}\hskip-13.00005pt+\hskip-15.3pt\underbrace{\mathbf{E}_{A,i}}_{\hskip 17.20844pt\text{residual}\in\mathbb{R}^{|A|}}\hskip-21.00009pt,$ (13)
where $\mathbf{E}_{A,i}\sim\mathcal{N}\bigl(0,(\boldsymbol{\Omega}^{\ast}_{A,A})^{-1}\bigr)$. Leveraging this relationship, the quantities in (12) can be further expressed as follows (Fan and Lv,, 2016):
@@
Appendix F.2 provides detailed derivations for (13) and (14). Since p|A|p-|A| can be comparable or larger than $n$, one may want to leverage the existing regularization method for fitting the regression model (13) to obtain sample estimates of $\boldsymbol{\Omega}^{\ast}_{A,A}$, $\mathbf{E}_{A,i}$ and $(\boldsymbol{\alpha}_{1})_{A}$, which further yields plug-in estimate of $\tilde{\mathbf{X}}_{A,i}$. This process can be repeated for all $A\in\mathcal{A}$ with $\mathcal{A}$ a partition set of $[p]$ consisting of sets of size 2. By combining estimates for $\tilde{\mathbf{X}}_{A,i}$ for all $A\in\mathcal{A}$, we can form an estimate of the full matrix $\tilde{\mathbf{X}}_{G_{1}^{*}}\in\mathbb{R}^{|G_{1}^{*}|\times p}$.
@@
Two issues remain to be addressed. First, the true cluster labels are unknown. Second, we need an implementable ISEE subroutine. Since our algorithm is iterative by nature, we use the cluster assignments in the current iteration as the proxy for true cluster labels to implement the ISEE subroutine, and then update the estimates for $\tilde{\mathbf{X}}$ and cluster assignments in the next iteration. We next explain the ISEE subroutine for a given pair of cluster assignment $(G_{1},G_{2})$.
ISEE Subroutine
The feature indices $[p]$ is partitioned into non-overlapping subsets $A_{1},\ldots,A_{L}$, each of size 2 if $p$ is even. For odd $p$, we set the size of $A_{L}$ to 3. The estimation procedure is then applied for each feature subset, and the results are aggregated. Without loss of generality, we describe the steps for the subset $A_{1}$. Given the previous cluster estimates $\hat{G}_{1}^{t-1}$ and $\hat{G}_{2}^{t-1}$, we treat $\{(\mathbf{X}_{A_{1}^{C},i},\mathbf{X}_{A_{1},i}):i\in\hat{G}_{1}^{t-1}\}$ as (predictor, response) pair and apply off-the-shelf high-dimensional regression method (for example, lasso; Tibshirani,, 1996) to estimate the intercept $(\boldsymbol{\alpha}_{1})_{A_{1}}$ and residuals $\{\mathbf{E}_{A_{1},i}\}_{i\in\hat{G}_{1}^{t-1}}$. Similarly, using $\{(\mathbf{X}_{A_{1}^{C},i},\mathbf{X}_{A_{1},i}:i\in\hat{G}_{2}^{t}\}$, we estimate the corresponding $(\boldsymbol{\alpha}_{2})_{A_{1}}$ and residuals $\{\mathbf{E}_{A_{1},i}\}_{i\in\hat{G}_{2}^{t-1}}$ for cluster 2. To estimate $\boldsymbol{\Omega}^{\ast}_{A_{1},A_{1}}$, we exploit the fact that the residuals $\mathbf{E}_{A_{1},i}\sim\mathcal{N}\bigl(0,(\boldsymbol{\Omega}^{\ast}_{A_{1},A_{1}})^{-1}\bigr)$, and use the inverse of the estimated sample covariance matrix of $\mathbf{E}_{A_{1},i}$ as an estimator. Then applying (14), we obtain the estimates for $(\tilde{\boldsymbol{\mu}}_{1}^{\ast})_{A_{1}}$, $(\tilde{\boldsymbol{\mu}}_{2}^{\ast})_{A_{1}}$, and $\{\tilde{\mathcal{E}}_{A_{1},i}\}_{i=1}^{n}$. Accordingly, using (12), we can estimate $\{\tilde{\mathbf{X}}_{A,i}\}_{i=1}^{n}$. Repeating this procedure for all subsets $A_{1},\ldots,A_{L}$ and stacking the results yields estimates of $\tilde{\boldsymbol{\mu}}_{1}^{\ast}$, $\tilde{\boldsymbol{\mu}}_{2}^{\ast}$ and $\tilde{\mathbf{X}}$. In summary, we have the output $\texttt{isee}(\mathbf{X},\hat{G}_{1}^{t-1},\hat{G}_{2}^{t-1})=(\hat{\tilde{\boldsymbol{\mu}}}_{1}^{t},\hat{\tilde{\boldsymbol{\mu}}}_{2}^{t},\hat{\tilde{\mathbf{X}}}^{t})$ in the $t$-th iteration. The full procedure is detailed in Algorithm 4 in Appendix F.2.
3.2.2 Full Algorithm with Iterative Feature Selection via ISEE
@@
Given the ISEE subroutine, Algorithm 3 extends Algorithm 2 to the unknown covariance scenario. In the feature selection step, we threshold $\hat{\tilde{\boldsymbol{\mu}}}_{1}^{t}-\hat{\tilde{\boldsymbol{\mu}}}_{2}^{t}$ obtained from ISEE subroutine, instead of $\bar{\tilde{\mathbf{X}}}_{\hat{G}_{1}^{t-1}}-\bar{\tilde{\mathbf{X}}}_{\hat{G}_{2}^{t-1}}$. In the clustering step, we replace $\tilde{\mathbf{X}}_{\hat{S}^{t},\cdot}$ with the estimated transformed data matrix $\hat{\tilde{\mathbf{X}}}^{t}_{\hat{S}^{t},\cdot}$. We replace $\boldsymbol{\Sigma}_{\hat{S}^{t},\hat{S}^{t}}$ with the pooled sample covariance, combining the sample covariances of $\mathbf{X}_{\hat{S}^{t},\hat{G}^{t-1}_{1}}$ and $\mathbf{X}_{\hat{S}^{t},\hat{G}^{t-1}_{2}}$. We leverage the convergence rate of the ISEE procedure (Fan and Lv,, 2016) to guide the choice of threshold for $\hat{\tilde{\boldsymbol{\mu}}}_{1}^{t}-\hat{\tilde{\boldsymbol{\mu}}}_{2}^{t}$. Specifically, for each $k\in\{1,2\}$ and for each subset $A_{\ell}$ with $\ell\in\{1,\ldots,L\}$, under some regularity conditions (Fan and Lv,, 2016), we have:
$\bigl\|\bigl((\hat{\tilde{\boldsymbol{\mu}}}_{1})_{A_{\ell}}-(\hat{\tilde{\boldsymbol{\mu}}}_{2})_{A_{\ell}}\bigr)-\bigl(({\tilde{\boldsymbol{\mu}}}_{1}^{\ast})_{A_{\ell}}-({\tilde{\boldsymbol{\mu}}}_{2}^{\ast})_{A_{\ell}}\bigr)\bigr\|_{2}\lesssim C_{6}\max\{(J\log p)/n,\sqrt{\log p/n}\},$ (15)
@@
where $J$ is the maximum number of nonzero off-diagonal entries per row of $\boldsymbol{\Omega}^{\ast}$, and $C_{6}=\|\boldsymbol{\Omega}^{\ast}_{A_{\ell},A_{\ell}}\|_{2}+\|(\hat{\boldsymbol{\alpha}}_{1})_{A_{\ell}}\|_{2}+\|(\hat{\boldsymbol{\alpha}}_{2})_{A_{\ell}}\|_{2}$. This bound comes from the convergence rates of $\|\hat{\boldsymbol{\Omega}}_{A_{\ell},A_{\ell}}-\boldsymbol{\Omega}^{\ast}_{A_{\ell},A_{\ell}}\|_{2}$ and $\|(\hat{\boldsymbol{\alpha}}_{k})_{A_{\ell}}-(\boldsymbol{\alpha}_{k})_{A_{\ell}}\|_{2}$ proved in Fan and Lv, (2016). The detailed derivation is provided in Appendix F.2. In the standard sparse high-dimensional regime of $J\sqrt{\log p/n}=o(1)$, the term $\sqrt{\log p/n}$ dominates the bound. Since ISEE only involves small submatrices of size $2\times 2$ or $3\times 3$, the terms in $C_{6}$ are bounded and contribute only constant factors. To account for these constants and avoid including too many noise features, we include a slowly diverging factor $\sqrt{\log n}$ in our threshold. Combining all the pieces, we set the feature selection threshold as $\sqrt{(\log n\log p)/n}$. Appendix H.2 shows numerically that this feature selection is slightly more conservative than the direct extension of (10).
@@
[…]
@@
4:   $\hat{\tilde{\boldsymbol{\mu}}}_{1}^{t},\hat{\tilde{\boldsymbol{\mu}}}_{2}^{t},\hat{\tilde{\mathbf{X}}}^{t}\leftarrow\texttt{isee}(\mathbf{X},\hat{G}_{1}^{t-1},\hat{G}_{2}^{t-1})\triangleright$ Algorithm 4
5:   $\hat{S}^{t}\leftarrow\left\{j\in[p]:|(\widehat{\tilde{\boldsymbol{\mu}}}_{1}^{t}-\widehat{\tilde{\boldsymbol{\mu}}}_{2}^{t})_{j}|>\sqrt{\log n\log p/n}\right\}$
@@
7:   $\hat{G}_{1}^{t},\hat{G}_{2}^{t}\leftarrow\texttt{clust}(\hat{\tilde{\mathbf{X}}}_{\hat{S}^{t},\cdot},\hat{\boldsymbol{\Sigma}}_{\hat{S}^{t},\hat{S}^{t}})$ $\triangleright$ Algorithm 1
@@
4 NUMERICAL STUDIES
For simulations, the data are generated from Gaussian distributions with two symmetric clusters:
@@ 4 NUMERICAL STUDIES
$\mathbf{X}_{i}\overset{i.i.d.}{\sim}\mathcal{N}(\boldsymbol{\mu}^{\ast},\boldsymbol{\Sigma}^{\ast}),~i=1,\dots,n/2,\quad\text{and}\quad\mathbf{X}_{i}\overset{i.i.d.}{\sim}\mathcal{N}(-\boldsymbol{\mu}^{\ast},\boldsymbol{\Sigma}^{\ast}),i=n/2+1,\dots,n.$ (16)
@@
We fix $S_{0}=[10]$ and increase the ambient dimension $p$. We assign equal magnitude to all entries of $\boldsymbol{\mu}^{\ast}_{S_{0}}$, with the magnitude chosen so that the signal strength, measured Mahalanobis distance $\Delta_{\boldsymbol{\Sigma}^{\ast}}=2\sqrt{(\boldsymbol{\mu}^{\ast})^{\top}\boldsymbol{\Omega}^{\ast}\boldsymbol{\mu}^{\ast}}=\sqrt{(\boldsymbol{\beta}^{\ast})^{\top}\boldsymbol{\Sigma}^{\ast}\boldsymbol{\beta}^{\ast}},$ remains constant as $p$ increases. We evaluate performance using the mis-clustering rate, averaged over 200 independent replications. We set the number of iteration as $T=100$, but used early stopping criterion. The details on this criterion, code implementation, and further information required for replicating our results are provided in Appendix G.1.
Table 1: Summary of Parameter Settings for (16). $i,j\in[p]$.
Covariance $p$ $n$ Separation
known ($\boldsymbol{{\Sigma}}^{\ast}=\mathbf{I}_{p}$) 50-5000 200 $\Delta^{2}=4^{2},5^{2}$
unknown ($\Omega^{*}_{i,i}=1$, $\Omega^{*}_{i,i+1}=\Omega^{*}_{i+1,i}=\rho$,
$\Omega^{*}_{i,j}=0$ o.w., $\rho\in\{0.2,0.45\}$ )
50-400 500 $\Delta_{\boldsymbol{\Sigma}^{\ast}}^{2}=3^{2},4^{2}$
Simulation under Known Covariance
@@
We evaluate Algorithm 2 under the setting of line 1 of Table 1. We first compare non-sparse clustering methods, including spectral clustering (Han et al.,, 2023), hierarchical clustering, and SDP-relaxed $K$-means with all features (Chen and Yang,, 2021), with Algorithm 2 initialized by each of them. Here, initialization refers to a rough cluster assignment $\hat{G}_{1}^{0},\hat{G}_{2}^{0}$ used to guide the first feature selection step. Results in Figures 1-(a) and 1-(b) show that while the non-sparse methods fail to adapt to sparsity, Algorithm 2, when properly initialized, is robust to high-dimensional noise and effectively adapts to sparsity. Even moderately good initializations are sufficient, as evidenced by the improved performance of Algorithm 2 when initialized with spectral or SDP $K$-means clustering. In contrast, poor initializations can hinder the clustering performance, as reflected in the weaker performance of Algorithm 2 initialized by hierarchical clustering. Next, Figures 1-(c) and 1-(d) highlight the competitiveness of our approach against existing two-step and iterative sparse clustering approaches: influential feature PCA (IFPCA; Jin and Wang,, 2016), sparse alternate similarity clustering (SAS; Arias-Castro and Pu,, 2017), sparse $K$-means (SKM; Witten and Tibshirani,, 2010) and CHIME (Cai et al.,, 2019). Details of these baseline methods are provided in Appendix G.2. Appendix H.1 presents results under non-isotropic covariances.
@@
Figure 1: Mis-clustering rates of Algorithm 2 compared to non-sparse and sparse baselines. The dotted lines (non-sparse methods) can also be interpreted as the mis-clustering rates of Algorithm 2 immediately after initialization, before any iteration. The settings are provided in line 1 of Table 1.
@@
Figure 2: Mis-clustering rates of Algorithm 3, compared to other sparsity-aware baseline methods under unknown covariance scenario with varying conditional correlations. The settings follow the line 2 of Table 1.
Simulation under Unknown Covariance
Algorithm 3 is evaluated in the setting with a chain-graph sparse precision matrix with conditional correlation $\rho\in\{0.2,0.4\}$, following the setup described in line 2 of Table 1. The results in Figure 2 show that our method outperforms other sparsity-aware clustering approaches. Appendix H.2 demonstrates the effectiveness of bypassing full precision matrix estimation. It also illustrates Algorithm 3 with an alternative feature selection threshold, demonstrating that an exact threshold is unnecessary and constant factors in thresholds are not a big concern. Appendix H.3 demonstrates the convergence of the iterative algorithm and shows that slight overestimation of $S_{0}$ does not affect clustering performance.
Application to Real Datasets
We apply Algorithm 2 to leukemia subtype clustering using gene expression data ($n=45$, $p=3571$; Golub et al.,, 1999) and Algorithm 3 to clustering pre-processed MNIST images of digits 1 and 7 ($n=1000$, $p=784$). The leukemia dataset, originally created by Golub et al., (1999), comprises 72 gene expression profiles of 3871 genes, obtained from bone marrow and peripheral blood samples of patients with two types of leukemia: acute myeloid leukemia (AML) and acute lymphoblastic leukemia (ALL). This dataset was previously analyzed by Jin and Wang, (2016), and we used the data file available from their repository. To assess statistical performance, we performed stratified subsampling of 45 samples, repeating this procedure 100 times to compute the average clustering accuracy. The MNIST dataset consists of $28\times 28$ grayscale images of digits from 1 to 10. We trained a $k$-sparse autoencoder (Makhzani and Frey,, 2013) on 30000 samples spanning all digits, using a sparsity parameter of $k=20$. The trained model was then applied to 8640 test images of digits ‘1’and ‘7’, converting them into 784-dimensional sparse representations. We applied stratified sub-sampling by selecting $n=1000$ images from the 8640 test samples, applied the clustering algorithms to distinguish digits ‘1’and ‘7’, repeated the process 100 times, and reported the average clustering accuracy. We compared Algorithms 2 and 3 with Sparse $K$-means (SKM; Witten and Tibshirani,, 2010), Influential Feature PCA (IFPCA; Jin and Wang,, 2016), and Sparse Alternate Similarity clustering (SAS; Arias-Castro and Pu,, 2017). We do not report the performance of CHIME (Cai et al.,, 2019) in this real-data analysis, as the original paper does not share the implemented code for initializing cluster centers and covariance, and unlike in simulations, there is no obvious choice of initialization for real datasets. Details of the baseline methods are provided in Appendix G.2. For the $k$-sparse autoencoder, we used the authors’ implementation available on their GitHub repository. The results in Table 2 show that our methods consistently outperform baseline approaches.
Table 2: Clustering accuracy on the leukemia gene expression dataset and the MNIST dataset.
Our method IFPCA SKM SAS
Leukemia 0.93 0.84 0.79 0.87
MNIST 0.94 0.61 0.57 0.56
Additional Simulations
Appendix H.4 demonstrates that our algorithms maintain strong clustering accuracy under mild violations of the assumptions of Gaussianity, covariance homogeneity, and exact sparsity of model parameters. For scalability, while our focus is on the high-dimensional, low-sample-size regime, real-world datasets may involve much larger samples. For larger $n$, we recommend approximate solvers (Zhuang et al.,, 2022, 2024) tailored for the SDP step. Appendix H.5 demonstrates that those approximate solvers scale our algorithms up to n=100,000n=100{,}000. Finally, we compare our methods to subspace clustering algorithms that extracts, rather than selects, features lying in low-dimensional subspaces (Soltanolkotabi et al.,, 2014; Wang and Xu,, 2016; Huang and Gu,, 2025). Specifically, Appendix H.6 shows that our method excels the Johnson–Lindenstrauss transform-based $K$-means (Boutsidis et al.,, 2010) in sparse settings.